### v1

# Fast-Revisit Coverage Path Planning for Autonomous Mobile Patrol Robots Using Long-Range Sensor Information

Abstract
The utilization of Unmanned Ground Vehicles (UGVs) for patrolling industrial sites has expanded significantly. These UGVs typically are equipped with perception systems, e.g., computer vision, with limited range due to sensor limitations or site topology. High-level control of the UGVs requires Coverage Path Planning (CPP) algorithms that navigate all relevant waypoints and promptly start the next cycle.
In this paper, we propose the novel Fast-Revisit Coverage Path Planning (FaRe-CPP) algorithm using a greedy heuristic approach to propose waypoints for maximum coverage area and a random search-based path optimization technique to obtain a path along the proposed waypoints with minimum revisit time. We evaluated the algorithm in a simulated environment using Gazebo and a camera-equipped TurtleBot3 against a number of existing algorithms. Compared to their average revisit times and path lengths, our FaRe-CPP algorithm approximately showed a 45% and 40% reduction, respectively, in these highly relevant performance indicators.
Index Terms:
Motion and Path Planning, Reactive and Sensor-Based Planning, Vision-Based Navigation, Surveillance Systems, Industrial RobotsI Introduction
Robots autonomously patrolling environments like industrial, manufacturing, construction or agricultural sites are supposed to conduct efficient surveillance by covering a maximum area in minimum time. In order to ensure this, a number of Coverage Path Planning (CPP) approaches exist that compute paths passing through all predefined or computed waypoints while covering the area of interest in the least time and with the least spatial effort. Robots rely on CPP to ensure efficient coverage, which is essential for optimal performance and safety in many applications. These algorithms can be categorized into online and offline approaches [1] where online approaches generally assume little to no prior knowledge about the environment and often use heuristics to ensure maximum coverage. Offline approaches utilize pre-recorded maps of the environment, making them well-suited for large-scale areas. By leveraging hardware capabilities, all these approaches can ensure an efficient path that maximizes area coverage. Furthermore, CPP algorithms are categorized based on various factors, including the environment type (indoor, outdoor), dimensionality (2D occupancy maps, 3D occupancy maps), and the platform (unmanned ground (UGV) or aerial (UAV) vehicles). Additionally, multi-robot CPP approaches exist [2] which are more efficient for large areas due to task division and faster coverage, but require complex coordination and may not be applicable in all scenarios.
Ensuring minimum revisit times of any given waypoint is crucial for certain application, for instance, safety patrol on hazardous substance sites like garbage dumps and recycling centers. Typically, in such sites, spontaneous fires are an imminent hazard because of inflammable substances or self-ignition because of the heat produced during composting. Therefore, fire incidents happen frequently [3, 4] and there is great interest by waste disposal enterprises to prevents such indicents. A safety patrol robot that is able to detect fires autonomously, and potentially even applies extinguishing agents directly, calls the fire department and keeps the fire at bay until their arrival. For such kind of patrols, minimum revisit times with maximum coverage are essential to prevent fires from getting out of control.
In this paper, we present a novel Coverage Path Planning algorithm with focus on fast revisit times for a single mobile patrol robot. We name the algorithm Fast-Revisit Coverage Path Planning (FaRe-CPP). Utilizing a pre-recorded 2D occupancy grid map, our approach updates the surveilled region within the field of view of the relevant onboard sensors at each step, subsequently creating waypoints that maximize the coverage area. Initially generated with a greedy heuristic for maximum perception potential, these waypoints are optimized using the GRASP method [5] to minimize path length while preserving coverage information. The combination of waypoint generation subsequent optimization ensures maximum overall area coverage, with no path overlap and minimal path length, resulting in minimal revisit times.
A Python implementation of FaRe-CPP with ROS111Robot Operating System, https://ros.org connectivity is publicly available on https://github.com/hcr-lab/fare-cpp
II Related Work
II-A Early foundations for Coverage Path Planning
Coverage Path Planning is essential for robot applications like vacuum cleaning [6], radiation monitoring [7], bridge inspection [8], emergency evacuation [3], and construction monitoring [9].
Earlier researchers implemented decomposition approaches for CPP, including exact cellular [10, 11], trapezoidal [12], Boustrophedon [13], and Morse-based [14] decomposition which divides free space into navigable cells, optimizing systematic path coverage through various cell partitioning and merging techniques. Miao et al. [15] propose a scalable CPP method for cleaning robots using rectangular map decomposition, simplifying the planning process for large-scale applications. These decomposition methods focus on the robot’s physical footprint rather than sensor ranges, such as camera fields of view, which are vital for scanning environments. They create zig-zag patterns and are less suited for patrolling robots, as they might disrupt ongoing activities. Gabriely et al. [16] discuss the Spanning Tree Covering algorithm, one of the first grid-based algorithms for CPP.
Latest methods in coverage path planning (CPP) move beyond merely partitioning areas and integrate both operational aspects and sensor capabilities into the area decomposition process [17, 18, 19, 8, 20, 21, 22, 23, 24]. For instance, Swain et al. [17] introduced an efficient path-planning algorithm for 2D ground area coverage using multi-UAVs, emphasizing the integration of sensor capabilities into the planning process. Mansouri et al. [18] developed a method for 2D visual area coverage and path planning that couples the robot’s path with the camera footprints, enhancing the fidelity of CPP by considering visual sensor constraints. Similarly, Ghaddar et al. [19] presented an energy-aware grid-based CPP for UAVs, incorporating area partitioning in the presence of no-fly zones, which highlights the importance of operational constraints in path planning. Phung et al. [8] proposed an enhanced discrete particle swarm optimization algorithm for UAV vision-based surface inspection, demonstrating the benefits of considering sensor footprints for efficient inspection tasks. Kantaros et al. [23] explored distributed coverage control for concave areas, considering visibility sensing constraints, further highlighting the need to integrate sensor capabilities into the CPP process. An et al. [25] introduced the Rainbow Coverage Path Planning (RCPP) method for patrolling mobile robots with circular sensing ranges, optimizing the path to minimize overlap in sensing regions and enhance patrolling efficiency.
II-B Advancements in Coverage Path Planning approaches
Tran et al. [22, 21, 24] contributed to the dynamic and robust multi-robot coverage of both known and unknown environments through frontier-led swarming techniques. These methods underscore the importance of dynamic adaptation and sensor integration in CPP. Baras and Dasygenis [20] focused on energy-efficient CPP for UGVs by reducing turns, crucial for prolonging operational durations and enhancing efficiency. Jonnarth et al. [26] proposed a deep reinforcement learning approach where the CPP problem is formulated as a Markov Decision Process. Talavera et al. [27] developed an autonomous ground robot to support firefighters in indoor emergencies. This robot utilizes advanced CPP algorithms to navigate complex indoor environments, detect fires, and provide real-time situational awareness to firefighting teams, demonstrating the practical utility of CPP in critical applications. Nasirian et al. [28] addressed the challenge of effectively covering and disinfecting an environment with mobile robots using a graph-based representation. Similarly, Pierson et al. [29] focused on deploying a mobile UVC disinfection robot with an optimized CPP algorithm. Perminov et al. [30] presented a genetic-based human-aware CPP algorithm for autonomous disinfection robots, achieving significant coverage efficiency. Chen et al. [31] utilize building information modeling (BIM) to enhance the efficiency of CPP for indoor robots, considering various robotic configurations for adaptable planning. This work builds on previous research [32] which explored global path planning based on BIM and physics engines for UGVs in indoor environments. Full coverage planning by Lee et al. [33] researches intricacies of path planning in environments with negative obstacles.
II-C Established practical implementations of Coverage Path Planning
Building upon the significant advancements in coverage path planning (CPP) techniques, certain studies have not only concentrated on theoretical models but have also ventured into practical implementations and analyses through software tools. The early work of Bormann et al. [34] an exploration of room segmentation proposed a set of suitable map division algorithms. Following this foundational work, they presented a comprehensive survey alongside the implementation and analysis of various indoor CPP methods [35].
III Proposed Approach: Fast-Revisit Coverage Path Planning (FaRe-CPP)
In this section, we discuss in detail the methodology of our novel Fast-Revisit Coverage Path Planning (FaRe-CPP) algorithm. Our main assumption is that the environment is represented as an occupancy map recorded using Simultaneous Localization and Mapping (SLAM) (see Section IV-A). The grid cells in the occupancy map fall into one of three categories: occupied, unoccupied, and unknown.
In FaRe-CPP, instead of considering the robot’s physical dimensions as the footprint, we utilize the field of view (FoV) of the range sensor, e.g. 2D camera, mounted on the robot as the footprint. The range within which the sensor can perceive range information is referred to as the sensor range. The grid cells within the FoV are then referred to as explored cells, and the center cell is considered the robot’s position (x, y) while the orientation () is obtained based on the direction of the robot relative to the coordinate system of the grid map. As visually summarized in Fig. 2, there are three main steps in FaRe-CPP which we describe in the next subsections.
III-A Step 1: Waypoint Candidate Generation
The algorithm begins with the waypoint generation, where the robot’s initial position is used as the first waypoint. The robot’s initial FoV is then calculated based on its orientation and sensor range. For each grid cell within the sensor range , the angle is determined using . If lies within the FoV range and is not blocked by an obstacle (checked using the Bresenham line algorithm [36]), the cell is marked as explored.
After updating the grid cells within the FoV to the explored region, the boundary cells that separate the explored region from the unoccupied region as shown in Fig. 3 serve as waypoint candidates . These waypoint candidates represent possible future positions for the robot from which the next actual waypoint is selected Step 2.
III-B Step 2: Coverage Area Estimation per Waypoint Candidate
After generating all waypoint candidates using the waypoint generation step, the coverage area is estimated for each waypoint candidate :
where and represent the cell indices in the grid map , represents the side length of each quadratic cell so that indicates the physical area of each cell, and represents the explored area of each sub-grid map with the respective waypoint candidate selected as next actual waypoint.
Separate sub grid maps per waypoint candidate are created, and the unoccupied region is updated to the explored region like in Fig. 4 as discussed in Step 1. The coverage area for each grid map with the respective candidate as next waypoint is computed.
III-C Step 3: Waypoint Selection by Maximum Coverage Area
After the coverage area has been computed for each waypoint candidate, the candidate with the potential of providing the maximum FoV is selected as the next waypoint with
The sub-grid map of this waypoint is then used in the next iteration where new waypoint candidates are generated.
The FaRe-CPP algorithm repeats Steps 1–3 until the stopping criterion is met. This is the case when no more waypoint candidates are available, a predefined minimum coverage area has been exceeded, or the rate of increase in the coverage area across iterations is less than a threshold :
III-D Step 4: Waypoint Optimization
So far, the FaRe-CPP algorithm focuses on incrementally expanding the coverage area by selecting waypoints that maximize the coverage area at each iteration, rather than minimizing the overall path length. As a result, the waypoints are not arranged in an optimal sequence, leading to inefficient, overlapping, or zigzagging paths. This disorganization is since the waypoints are obtained in The waypoint generation step purely based on coverage needs, not on path efficiency. To balance extensive coverage with path efficiency, we employ the Greedy Randomized Adaptive Search Procedure (GRASP) [5]. GRASP operates in two phases: 1.) initial greedy randomized construction, which builds a path by selecting from top candidates, introducing randomness to avoid local optima, followed by 2.) local search, which refines the path using the 2-opt swap strategy to minimize cumulative distance.
To avoid computational complexity, the environment map is also not considered during waypoint optimization. The optimization focuses only on ordering the waypoints, while path planning that takes the environment into account is done in the next step. GRASP optimizes the waypoints such that the global path that passes through the waypoints bears the least cumulative distance, as shown in Fig. 5. However, due to the introduction of randomness in the greedy construction phase and the iterative nature of the 2-opt swap strategy, there is some variability in the resulting global path. While the global path is optimized, the exact distance can vary slightly across different runs due to the stochastic elements of the algorithm. Despite this, GRASP consistently arranges waypoints with reduced cumulative distance and improved efficiency.
III-E Path Planning and Following
After generating and optimizing waypoints, they are used to plan a global path. In this case, the A* algorithm [37] is employed to find a path that passes through these waypoints. A* is a graph-based search algorithm that finds the shortest path from a start point to a goal, considering the cost of movement between points. The result is a global path that connects the waypoints while maintaining alignment with the range sensor’s field of view. This path, which forms a closed loop through all waypoints, can be utilized by and path following approach to locomote the robot accordingly. In the implementation of our example scenario as used in the following experimental evaluation, we use the Dynamic Window Approach (DWA) from the ROS navigation stack for local path planning.
IV Evaluation
IV-A Experimental setup
We conducted all experiments in a simulation environment using ROS Noetic, Gazebo, and a camera-equipped TurtleBot3. We first created an occupancy grid map using the SLAM technique with the GMapping package [38]. We selected two environments from AWS RoboMaker: The House world222https://github.com/aws-robotics/aws-robomaker-small-house-world for comparison with algorithms designed for smaller environments and the Warehouse world333https://github.com/belal-ibrahim/dynamic_logistics_warehouse for larger environments like warehouses or recycling centers.
The occupancy grid map of the House environment has dimensions of 500x500 grid cells with a resolution of , resulting in an unoccupied coverage area of , whereas the Warehouse environment has 380x640 grid cells with a resolution of and an unoccupied coverage area of .
IV-B Evaluation Metrics
We evaluated the performance of FaRe-CPP on the generated grid map using several key metrics, chosen to comprehensively assess the algorithm’s efficiency and effectiveness in selected environments:
Path Length : This metric measures the total distance traveled by the robot in a closed loop, starting and ending at the first waypoint. It provides insight into the efficiency of the generated path, with shorter paths being more optimal with respect to minimum revisit times.
Total Cumulative Rotation : This is calculated based on the robot’s orientation changes at each waypoint. While maintaining maximum coverage, this should be minimized in order to achieve fast revisit times.
Revisit Time : Revisit time measures the duration required for the robot to complete its patrol loop and return to the starting point for subsequent patrols. It is determined as with as the actual linear and as the actual angular robot velocity.
Coverage Area: Percentage of the environment’s floor plan covered by the sensor’s field of view.
Computation Time: Total time to compute the final resulting path.
IV-C Results
As range sensors, we employ wide-angle cameras with FoVs of and on the Warehouse map, each with a 5-meter sensor radius. We use this radius as a conservative estimate to prove our algorithm’s utility because typical low-cost sensors as used on patrol robots have a range. Table I shows that the FoV outperforms the FoV, with shorter computation time ( vs. ), better coverage ( vs. ), and reduced cumulative rotation. These results demonstrate the effectiveness of FaRe-CPP in large environments like the Warehouse, see Fig. 6. The algorithm achieves high coverage and minimizes cumulative rotation, contributing to efficient sensor usage. Importantly, the similar revisit times ( for FoV and for FoV) indicate that FaRe-CPP maintains consistent performance without unnecessary re-exploration, making it well-suited for large-scale applications where extensive coverage and minimized redundancy are critical.
| Metric | FoV() | FoV() |
| Path Length (m) | 100.95 | 102.15 |
| Total Cumulative Rotation (rad) | 43.98 | 40.84 |
| Revisit Time (s) | 421 | 419 |
| Coverage Area (%) | 94 | 96 |
| Computation Time (s) | 178 | 149 |
IV-D Comparison with Existing Algorithms
| Algorithm | Computation Time (s) | Path Length (m) | Total Rotation (rad) | Coverage Area (%) | Revisit Time (s) |
| Grid TSP[35] | 148 | 297.45 | 544 | 97 | 2037 |
| Boustrophedon[13] | 4 | 209.79 | 302 | 76 | 1280 |
| Neural Network[39] | 8 | 293.69 | 795 | 92 | 2507 |
| Convex SPP[40] | 17 | 171.60 | 347 | 92 | 1239 |
| Grid Local Energy[41] | 1.7 | 261.34 | 322 | 95 | 1490 |
| Contour Line[35] | 71 | 283.74 | 574 | 97 | 2049 |
| Proposed FaRe-CPP | 109 | 148.00 1.66 | 227 0 | 95 0 | 948 5.53 |
In this section, we demonstrate our proposed FaRe-CPP’s effectiveness by comparing it to other algorithms that utilize camera footprints qualitatively in Fig. 7 and quantitatively in Table II. As a baseline, we compare our work to the algorithms thoroughly investigated in Bormann et al.’s survey [35] in their well-established, freely available implementations444https://github.com/ipa320/ipa_coverage_planning. We applied these algorithms to our chosen environment grid map, maintaining consistency with their work by using the same parameters for the field of view. Specifically, we considered an RGB-D camera with a horizontal field of view of at the robot’s position and a sensor radius of . The revisit time is determined as explained above with and .
Based on the results shown in Fig. 7, our proposed FaRe-CPP algorithm achieves more efficient coverage with fewer overlaps and a more streamlined path compared to the other methods. Table II quantitatively supports these findings, where we present averages for 10 comparison runs. FaRe-CPP exhibits the shortest path length, lowest total cumulative rotation, and fastest revisit time while maintaining a high coverage area percentage. FaRe-CPP showing no variance in computation time, total rotation and coverage area is based on the fact that these metrics are calculated in the the heuristic, deterministic parts of the algorithm. In total, these metrics highlight the performance of FaRe-CPP in achieving near-optimal coverage with fast revisit times for subsequent patrolling, which is an essential requirement for the targeted use cases.
With respect to computing time, FaRe-CPP currently performs worse than the other algorithms in their respectively used implementation except for Grid TSP. The main reason is that the currently available implementation of our approach uses Python where all others are implemented in C++. While FaRe-CPP works heuristically and often finds near-optimal solutions, its computation time can be exponential, especially as the size and complexity of the grid map increase.
FaRe-CPP performs well in coverage efficiency as well as path optimization and the main issue of the application problem at hand is closed-loop revisit time after initial deployment. Since no near-realtime reprocessing of the path is required after the robot started to patrol, any reprocessing can happen during deployment while the robot is on patrol. Initial computation, in contrast, plays a comparably minor role since the main focus of our approach is to achieve fast revisit times.
We provide the experimental results described in this section as generated with the current version of FaRe-CPP together with the code on https://github.com/hcr-lab/FaRe-CPP/tree/main/experiments.
V Conclusion
The proposed FaRe-CPP coverage path planning algorithm achieves near-complete coverage of the environment with short paths and low total cumulative rotation, resulting in fast revisit times. Our approach effectively first generates waypoints that yield the maximum visual coverage, and subsequently optimizes the waypoints using the Greedy Randomized Adaptive Search Procedure (GRASP).
We evaluated our approach in a simulation environment, demonstrating that our algorithm outperforms existing approaches by achieving optimal coverage areas with shorter path lengths, lower cumulative rotation, and faster revisit times. Compared to the average revisit times and path lengths of existing approaches, FaRe-CPP approximately showed a 45% and 40% reduction, respectively, in these highly relevant performance indicators.
However, so far FaRe-CPP is limited to scenarios with static obstacles and consistent coverage parameters since the path is being generated offline. In future work, we aim to expand FaRe-CPP to be able to deal with dynamic environments with online obstacle avoidance and varying coverage requirements. This will help to further improve the robustness and adaptability of FaRe-CPP in real-world scenarios.